### v2

# Variational Methods for Computing Non-Local Quantum Strategies

Abstract
In a nonlocal game, two noncommunicating players cooperate to convince a referee that they possess a strategy that does not violate the rules of the game. Quantum strategies allow players to optimally win some games by performing joint measurements on a shared entangled state, but computing these strategies can be challenging. We develop a variational algorithm for computing strategies of nonlocal games and show that it can yield optimal strategies for small examples of both convex and non-convex games. We show that our algorithm is capable of generating a short-depth circuit that implements an optimal quantum strategy for a graph coloring game. Moreover, we describe how this technique can be run on quantum computers and argue that such circuits will be useful for benchmarking because of their sensitivity to 2-qubit gate noise and application to self-testing. Finally, we demonstrate the use of these strategies experimentally on 11 IBM quantum computers.
1 Introduction
Quantum non-local games provide a unique setting for exploring the relationship between classical physics, quantum theory, and other non-signaling theories [1, 2, 3, 4]. An extensive body of research links these games to foundational problems in quantum physics, conjectures in operator algebras, and computational complexity theory. [5, 6, 7]. Moreover, advances in quantum information theory and combinatorics have revealed broad classes of games with a probable quantum advantage when players are allowed to incorporate quantum resources into their strategies [8, 9], making them exciting experimental candidates for self-testing and benchmarking of quantum hardware [10, 11].
However, despite our theoretical understanding of these games, constructing optimal strategies for nonlocal games remains a challenge. We propose a new methodology to construct strategies using variational algorithms to address this challenge.
1.1 Setup
A nonlocal game of $N$ players $\mathcal{G}=(Q_{1},...,Q_{N},A_{1},...,A_{N},\lambda)$ (illustrated in Fig. 1) consists of a set of possible questions $Q_{i}$ that player $i$ receives from a referee and a set of answers $A_{i}$ that player $i$ is allowed to send to the referee, which the referee then evaluates against a rule function $\lambda:Q_{1}\times...\times Q_{N}\times A_{1}\times...\times A_{N}\to\{0,1\}$. The game proceeds in the following steps:
  1. 1.
    The players are informed about the rules of the game $\lambda$, and can collaborate to create a joint strategy, modeled as a conditional probability density between the questions and answers, to maximize their chances of satisfying the rules of the game before it starts.
  2. 2.
    Players are then separated or isolated to prevent them from communicating. This is referred to as non-signaling, or in other words, each player’s actions are independent of each other.
  3. 3.
    The referee tests the strategy by asking questions to each player $\mathbf{q}=[q_{1},...,q_{n}]$ and receiving their responses $\mathbf{a}=[a_{1},...,a_{n}]$, where $q_{i}\in Q_{i}$ and $a_{i}\in A_{i}$, respectively.
  4. 4.
    The players win a round if $\lambda(\mathbf{a}|\mathbf{q})=1$, and lose if $\lambda(\mathbf{a}|\mathbf{q})=0$. Multiple rounds are played with different questions as required for the referee to be confident in the accuracy of the strategy.
Figure 1: Flow of a nonlocal game. After formulating a strategy, Alice and Bob separate and cannot communicate. For a quantum strategy, they each take a part of an entangled state $\rho$ and upon receiving a question $q$ from the referee, they perform a measurement on their respective states, giving an answer $a\sim p(a|q)$. Finally, the referee receives their answers and verifies them against the rules $\lambda(a|q)$.
Games are frequently synchronous, where all players share the same set of possible questions $Q$ and answers $A$. Synchronous games have rules that require the answers of two (or more) players to be identical when asked the same questions $\lambda(a_{i},a_{j},\dots|q,q,\dots)=0,~{}\forall a_{i}\neq a_{j}$. In our work, we only consider computing strategies for synchronous games.
Using the rules, we can define the value of the game as
$V(\mathcal{G})=\sum_{qa}\lambda(a|q)p(q)p(a|q),$ (1)
where the sum is taken over all possible values of $q\in Q^{N}$ and $a\in A^{N}$ (we drop the vector notation for convenience). The distribution $p(q)$ of the questions asked is typically chosen to be uniform, and the behavior $p(a|q)$ is determined by the strategy that the players construct. Notice that this is the only term that players can control to maximize their win rate. A strategy is said to be perfect if $\lambda(a|q)=0\implies p(a|q)=0$ and, consequently, $V(\mathcal{G})=1$.
Classical strategies consist of a lookup table that indexes each player’s response to a particular question. It suffices to consider deterministic strategies since stochastic ones cannot outperform deterministic strategies due to the linearity of the value of the game [3].
Suppose that players share a quantum state |$\ket{\psi}\in\otimes_{i}\mathcal{H}_{i}$, and each player has a set of POVMs with POVM elements of the form $\{\mathcal{M}_{a|q}\}$ that they perform on their subspace. Using this setup, players can generate the following conditional probability density,
$p(a|q)=\mathrm{Tr}\left[\rho\left(\otimes_{i}\mathcal{M}_{a_{i}|q_{i}}\right)\right].$ (2)
These densities are known as Quantum Strategies. We show in the Appendix that quantum strategies of this form can also be viewed as a quantum lookup table.
In addition to the above definition of a quantum strategy, there are a variety of competing definitions for a quantum strategy depending on the choice of axioms that describe how joint measurements between two parties should be performed [12]. In our work, we will only consider strategies as defined above, but the study of quantum strategies is a very active area of research [13, 14, 15, 16].
A game exhibits a quantum advantage if there exists a quantum strategy that performs better than the best possible classical strategy, in which case there is an inequality $\mathcal{I}$ that is violated for some quantum strategies. The inequality has historically served as an experimental test of local realism [17].
2 Method
In this section, we present a variational algorithm for computing quantum strategies in non-local games. Let |$\ket{\psi}$ be the shared entangled state between the players and $\mathcal{M}_{a|q}=\bigotimes_{i}\mathcal{M}_{a_{i}|q_{i}}$ be the joint POVM applied to that state for question $q$, returning $a$ with probability $p(a|q)=\langle\mathcal{M}_{a|q}\rangle$. It was noted by [18] that fixing these measurement operators gives a Hamiltonian whose ground state is the optimal shared state for those measurement settings. This fact has been used with reinforcement learning to optimize measurement settings while selecting the optimal shared state through exact diagonalization [19].
2.1 Dual-Phase Optimization
We build on this idea for our algorithm, which takes a VQE approach instead of diagonalizing. We call it dual-phase optimization (DPO), as it alternates between 2 phases: preparing the optimal state |$\ket{\psi}$ for the fixed measurements $\{\mathcal{M}_{a|q}\}$, and optimizing the measurements, fixing the shared state. We assume that the players parameterize their state |$\ket{\psi}\rightarrow\ket{\psi(\theta)}$ and POVMs $\mathcal{M}_{a|q}\rightarrow\mathcal{M}_{a|q}(\phi)$.
Algorithm 1 Dual-Phase Optimization
Initialize $\phi$ randomly
while $\Delta\langle H\rangle>\epsilon$ do
     Construct $H(\phi)$
     |$\ket{\psi(\theta)}\leftarrow VQE(H(\phi))$
     $\phi\leftarrow GD(\langle H(\phi)\rangle)$
end while
The preparation of the Hamiltonian depends on the specific measurement scheme the players decide on. Later, we outline a method for constructing a Hamiltonian from arbitrary game rules $\lambda$ and measurements $\{\mathcal{M}_{a|q}\}$.
The optimal shared state is prepared in the first phase using any VQE procedure $VQE(\cdot)$. Here, we choose ADAPT-VQE [20] because it generates compact variational circuits for use on near-term quantum hardware, but any other solver can also be used. The reference state |$\ket{\psi_{0}}$ can be a classical (product) state, e.g. |$\ket{0}$, $\ket{+}$. We choose a qubit operator pool consisting of all possible Pauli strings $P$ acting on the entire system. The operators added to the state take the form $e^{i\theta P}\rightarrow\ket{\psi(\theta)}=\prod_{j=N}^{1}e^{i\theta_{j}P_{j}}\ket{\psi_{0}}$, and they are capable of generating the entanglement required to win non-local games, provided they act non-trivially on at least 2 qubits.
The second phase uses a gradient descent-like optimizer $GD(\cdot)$ to update $\phi$. This requires the calculation of gradients $\nabla_{\phi}\braket{\psi(\theta)}{H(\phi)}{\psi(\theta)}$ on the quantum device, which can be done through parameter shift rules [21, 22].
2.2 Gradient Sample Complexity
The complexity of such a gradient calculation, in terms of the number of exponentials required to achieve a variance of $\epsilon_{0}^{2}$ for each component of the gradient is $\mathcal{O}(1/\epsilon_{0}^{2})$. To see this, let $\mathcal{E}_{j}$ be a random variable describing the error for experiment $j$. We then have $\mathbb{E}(\sum_{j}\mathcal{E}_{j}^{2})=N(\mathbb{E}(\mathcal{E}_{j}^{2}))=N\epsilon_{0}^{2}$. If we aim to ensure that we achieve standard deviation $\epsilon^{2}$ in the Euclidean norm squared, then it suffices to choose $\epsilon_{0}=\epsilon/\sqrt{N}$ and thus we can compute the expectation value within error $\epsilon$. As this has to be repeated $N$ times using our protocol, the number of operator exponentials needed to provide such an estimate of the gradient is
$N_{exp}\in\mathcal{O}\left(\frac{N^{2}}{\epsilon^{2}}\right).$ (3)
If we demand that the error per epoch of gradient descent is at most
$\gamma\min_{\theta\in\Gamma}\|\nabla_{\phi}\braket{\psi(\theta)}{H(\phi)}{\psi(\theta)}\|$, where $\Gamma$ is the set of parameters such that the parameters attained from the gradient descent is a subset of $\Gamma$, the total number of exponentials needed to ensure the above variance bound is
$N_{exp,tot}\in\mathcal{O}\left(\frac{N^{2}N_{\rm epoch}}{\gamma^{2}\min_{\theta\in\Gamma}\|\nabla_{\phi}\braket{\psi(\theta)}{H(\phi)}{\psi(\theta)}\|^{2}}\right).$ (4)
In general, the number of epochs is difficult to compute. In cases where the objective function is strongly convex, it can be shown to vary poly-logarithmically with the desired tolerance. Additionally, if there are barren plateaus, then almost all paths in $\Gamma$ will contain points with exponentially small gradients. Thus, we do not expect this variational algorithm to be generically efficient.
Since the parameters $\phi$ are randomly initialized, multiple runs of the algorithm may be required to find the optimal strategy. For the purpose of testing Bell inequalities, the best trial provides a bound on the optimal quantum value. The possibility of multiple trials is common in other approaches, such as the episodic nature of RL required for exploration [19][23].
2.3 Game Hamiltonians
As mentioned above, DPO requires the construction of a Hamiltonian based on the measurements of the players. Player $i$ may measure their qubits |$\ket{\psi}_{i}$ in an arbitrary basis depending on the question, leading to a form for the measurement operators
$\mathcal{M}_{a|q}$ $=\bigotimes_{i}U_{iq_{i}}^{\dagger}P_{a_{i}}U_{iq_{i}}$ (5)
$=U_{q}^{\dagger}P_{a}U_{q},$ (6)
where $P_{a_{i}}=\outerproduct{a_{i}}{a_{i}}$ | is the projector onto answer $a_{i}$, and $U_{iq_{i}}$ acts on |$\ket{\psi}_{i}$ in response to question $q_{i}$. Because $\langle\mathcal{M}_{a|q}\rangle=p(a|q)$, we can substitute this into (1) to construct an operator
$\beta$ $=\sum_{qa}\lambda(a|q)p(q)\mathcal{M}_{a|q}$ (7)
$=\sum_{q}p(q)U_{q}^{\dagger}\left(\sum_{a}\lambda(a|q)P_{a}\right)U_{q}$ (8)
with the property $\braket{\beta}=V(\mathcal{G})$. VQE finds the ground state of a Hamiltonian, so to maximize the win rate, we use a value Hamiltonian $H=-\beta$ in DPO. The problem can equivalently be framed as minimizing the error of the strategy with a substitution $\lambda\rightarrow 1-\lambda$, leading to a violation Hamiltonian $H=1-\beta$. This violation Hamiltonian is conceptually similar to loss functions in supervised machine learning.
Proposition 2.1.
The value $\braket{\beta}=1$ if and only if the players have a perfect quantum strategy, otherwise $\braket{\beta}<1$.
Proof 2.1.
We show this by first computing the value for a perfect strategy and then for an imperfect strategy. Let $I=\{(q,a)~{}|~{}\forall q,a~{}\lambda(a|q)=0\}$ be the set of question-answer pairs for which the strategy violates the rules, and let $P=I^{c}$ be its complement, the set of correctly answered question-answer pairs.
For a perfect quantum strategy, $I=\emptyset$ and $P=Q^{N}\times A^{N}$, therefore we get
$\braket{\beta}$ $=\sum_{q,a\in P\cup I}p(q)\lambda(a|q)$ $=\sum_{qa\in P}p(q,a)+(0)\sum_{qa\in I}p(q,a).$
Since $I=\emptyset$ for all $q,a$ pairs for which $\lambda=1$ are contained in $P$, and so the joint probability density in the left term must sum to 1. We obtain $\braket{\beta}=1$.
A very similar line of reasoning holds for an imperfect strategy, where $I\neq\emptyset$. Reusing the above expression,
$\braket{\beta}$ $=\sum_{qa\in P}p(q,a)+(0)\sum_{qa\in I}p(a,q)$
$=\sum_{qa\in P}p(q,a)<1,$
since $p(q,a),~{}q,a\in P$ no longer contains the full probability density as $I$ contains some possible pairs. We conclude that $\braket{\beta}\leq 1$, with $\braket{\beta}=1$ iff a strategy is perfect.
$\square$
To parameterize this Hamiltonian for DPO, a general single-qubit unitary $U_{1}$ may be decomposed into 3 parameters, leading to a parameterization of the full measurement gate $U_{q}=\bigotimes_{i,j_{i}}U_{1}(\phi_{iq_{i}j_{i}})$, where $i$ indexes the player, $q_{i}$ indexes the question, and $j_{i}$ indexes the particular qubit of player $i$. In measurement layers acting on multiple qubits, we expand each entry of $\phi_{iq_{i}j_{i}}$ to be the concatenated vector of parameters for all gates applied to that qubit, i.e. $U_{q}=\bigotimes_{i}U_{N_{i}}(\phi_{iq_{i}})$, where $U_{N_{i}}$ is an $N_{i}$-qubit unitary.
3 Experiments
To evaluate the performance of DPO, we apply it to two nonlocal games with known quantum bounds: CHSH and the N-partite symmetric (NPS) game [19]. Then, we use DPO to explicitly compute an optimal quantum strategy for the coloring game of a $14$ vertex graph called $G_{14}$ [24], for which we believe no strategy has been explicitly constructed. This strategy is then evaluated on real quantum hardware, demonstrating that it can be used as a benchmark for non-locality capabilities of quantum devices.
3.1 CHSH
The Clauser-Horne-Shimony-Holt (CHSH) scenario [17] is the simplest nonlocal game that admits a quantum advantage. CHSH features 2 players, Alice and Bob, who each receive a question $q_{i}\in Q=\{0,1\}$, answering $a_{i}\in A=\{0,1\}$. The inequality operator can be expressed in the familiar form $\mathcal{I}=A_{0}B_{0}+A_{1}B_{0}+A_{0}B_{1}-A_{1}B_{1}$ following the rules
$\lambda(a|q)=\begin{cases}a_{a}\oplus a_{b},&\text{if }q_{a}=q_{b}=1\\ \delta_{a_{a},a_{b}},&\text{otherwise}\end{cases}$ (9)
and making the substitution $\lambda=0\rightarrow\lambda=-1$ in (8). Here $A_{q}$ denotes Alice’s measurement operator and likewise $B_{q}$ for Bob. All classical strategies are bounded by $\braket{\mathcal{I}}\leq 2$, whereas quantum strategies can violate this up to $\braket{\mathcal{I}}=2\sqrt{2}$. It suffices to share a Bell state and then perform the appropriate single-qubit measurements.
We applied DPO to the CHSH game, which required only one trial, since it is a convex optimization problem [24]. $R_{y}(\phi)=e^{-i\frac{\phi}{2}Y}$ gates were chosen as the $U_{q}$ operators. In the first iteration, ADAPT chose |$\ket{\psi(\theta)}=e^{i\frac{\pi}{4}YX}\ket{00}$, which correctly generates a Bell state |$\ket{\Phi^{-}}$. In the second phase of that iteration, the measurement parameters were optimized to $\phi\approx[0,-\pi/2,\pi/4,-\pi/4]$ by constraining $\phi_{a0}=0$, giving the optimal inequality value $2\sqrt{2}$.
3.2 N-partite Symmetric
The NPS scenario [25] involves the correlations between players $N$, each receiving a binary question $q_{i}\in\{0,1\}$ and returning a dichotomic answer $a_{i}=\pm 1$. The inequality is expressed in terms of one- and symmetric two-body correlators,
$S_{q}=\sum_{i}\braket{\mathcal{M}_{iq}},~{}S_{qq^{\prime}}=\sum_{i\neq j}\braket{\mathcal{M}_{iq}\mathcal{M}_{jq^{\prime}}},$ (10)
where measurement $\mathcal{M}_{iq}=U_{iq}^{\dagger}Z_{i}U_{iq}$. The classical bound on the correlations is
$\mathcal{I}=-2S_{0}+\frac{1}{2}S_{00}-S_{01}+\frac{1}{2}S_{11}+2N\geq 0,$ (11)
with negative values only achievable with quantum strategies [19].
We tested 50 DPO trials for the $N=6$ case (Fig. 2). Our algorithm encounters some local minima, particularly at the classical bound of $\mathcal{I}=0$, but still succeeds in 19/50 attempts reaching $\mathcal{I}=-0.258$ as found in [19] as well. Additionally, 29/50 of the trials violated the classical bound.
Figure 2: Trials of DPO on NPS for $N=6$. (Left) Trajectory of all 50 trials. Negative inequality values are not reachable with classical states. (Right) Distribution of the final inequality values. Despite the non-convexity of the problem, many trials still reach the optimal value.
3.3 Chromatic Number Game
The objective of the chromatic number game [24] is to color a graph $G=(V,E)$ in such a way that adjacent vertices are never given the same color. This differs from the other nonlocal games we mentioned, as the sets of questions and answers are much larger, and each player requires more qubits to encode their answer. The referee asks a question $q=[v_{a},v_{b}]\in V\cup E$, and the players respond with colors $a=[c_{a},c_{b}],~{}c_{i}\in[c]$. The rules are given by
$\lambda(a|q)=\begin{cases}\delta_{c_{a},c_{b}}&\text{if }v_{a}=v_{b}\\ (1-\delta_{c_{a},c_{b}})&\text{if }v_{a}\neq v_{b}\end{cases}.$ (12)
The graph $G_{14}$ in Fig. 3 has been shown by [24] to have a quantum chromatic number $\chi_{q}(G_{14})=4$, while the classical chromatic number $\chi(G_{14})=5$, which admits a quantum advantage. However, a quantum strategy to achieve this has not yet been constructed. It should be noted that there is a way to construct a strategy using an orthogonal representation of $G_{14}$ in $\mathbb{R}^{4}$, outlined in [24], but it is not clear how to obtain an explicit set of ansätze from the representation to construct a short-depth circuit for the strategy. We use the DPO algorithm to generate a perfect quantum strategy that achieves $\chi_{q}(G_{14})=4$.
Figure 3: Picture of the $G_{13}$ graph. To get the graph $G_{14}$ just add an apex vertex $\alpha$ to this graph. Borrowed from [24].
For convenience, we denote a vertex question $[v,v]$ as $v$, and an edge question $[v_{a},v_{b}]\in E$ as $e$. Let the answers also be given by $c=[c_{1},c_{2}]$. Evaluating the expression in (8) gives
$\beta$ $=\frac{1}{|Q|}\left[\sum_{v}U_{v}^{\dagger}P_{cc}U_{v}+\sum_{e}U_{e}^{\dagger}(I-P_{cc})U_{v}\right]$ (13)
$=\frac{1}{|Q|}\left[\sum_{v}U_{v}^{\dagger}P_{cc}U_{v}-\sum_{e}U_{e}^{\dagger}P_{cc}U_{v}+|E|\right],$ (14)
where $P_{cc}=\sum_{c}\ket{cc}\bra{cc}$ | is the projector onto the space of equal colors (i.e. $p(c_{1}=c_{2}|q)=\braket{U_{q}^{\dagger}P_{cc}U_{q}}$), and $|Q|=|V|+|E|$. Intuitively, the first term maximizes $p(c_{1}=c_{2}|v)$, and the second term maximizes $p(c_{1}\neq c_{2}|e)=1-p(c_{1}=c_{2}|e)$. Note that to ensure that all possible questions are asked to each player, $E$ contains both edges $(v_{1},v_{2})$ and their reverse $(v_{2},v_{1})$.
To further simplify the search for strategies, we restrict the players to 2 qubits each, since 2 qubits suffice to represent $c\in\{0,...,\chi_{q}-1\}$. Additionally, [24] imposes constraints on the optimal quantum strategy, i.e., that Bob’s measurement operators are complex conjugate to Alice’s, halving the number of measurement parameters $\phi$ required. We use a measurement layer per player of general single-qubit $U$ gates, a CNOT from qubit 0 to qubit 1, and finally $R_{y}$ gates applied to each qubit, resulting in 8 parameters per question or 112 in total (in our code, this is the U3Ry layer).
We simulated 500 trials of DPO, achieving a minimum energy of $E=-1.0000$. We remove the gates added by ADAPT with $|\theta_{i}|<10^{-4}$. The corresponding circuit was then converted into a Qiskit circuit (Fig. 4), and the evaluation using the classical AerSimulator confirmed a 100% win rate, providing experimental evidence of the proof in [24]. For brevity, we exclude the 112 parameters $\phi$ in this paper, but they can be found in our code repository. The strategy was then submitted to 11 IBM quantum devices with 4 or more qubits (Fig. 5).
Figure 4: Generated circuit for $G_{14}$. The initial state |$\ket{+}^{\otimes N}$ is prepared, then ADAPT added the operators $Y_{0}Z_{2}$ and $Y_{0}Z_{1}Y_{2}Y_{3}$, giving the shared state |$\ket{\psi(\theta)}$. The remaining gate layers along with the 112 measurement parameters $\phi$ constitute the measurement strategy.
(a) Qiskit AerSimulator
(b) IBMQ ibm_hanoi
Figure 5: Win rate of the $G_{14}$ strategy by question. (Left) Classical statevector simulation with 10000 shots per question. (Right) Performance on a quantum device with 1024 shots per question.
The operators chosen by ADAPT are nonlocal as expected. The shared state
|$\ket{\psi(\mathbf{\theta})}$ $=e^{i\theta_{1}YZYY}e^{i\theta_{0}YIZI}\ket{+}$ (15)
$=\frac{1}{2}H^{\otimes 4}\sum_{c}\ket{cc}$ (16)
has a mutual information of $4$ between Alice and Bob, demonstrating the high degree of entanglement required to win non-local games. The discovered strategy exploits the fact that the referee cannot cross-check answers from previous questions (otherwise the graph would not be 4-quantum colorable), and the players change their coloring for each vertex probabilistically in subsequent runs, using the entanglement to coordinate their answers as required. For example, when asked $q=[A,A]$ multiple times, the responses are nearly uniform among $[c]$ but always match. Furthermore, we found that measurement layers consisting of only single-qubit gates were insufficient and generated imperfect strategies at $E=-0.9921$. In these cases, we frequently observed that the errors consisted of a cyclic path through some graph edges.
(a) All devices
(b) Noise Simulation
Figure 6: (Left) Average performance of the strategy on quantum devices grouped by question category, either a vertex $q=[v,v]$ or an edge $q=[v_{1},v_{2}]$ (vertex winrate on the left, edge winrate on the right). The number of device qubits is reported to distinguish processor types; the circuit was executed on 4 qubits. (Right) Classical simulation of the circuit with random Pauli noise applied to CX gates with probability $p_{err}$. The circuit was transpiled to the basis gates and coupling map of ibm_hanoi, and the observed data are fitted to the curves by maximizing the probability of observing the data assuming a normal distribution. The estimated error for CX gates is higher than reported, since our simulation did not account for measurement readout or decoherence errors.
3.4 Simulations on IBM Devices
We saw a decrease in performance on IBM quantum devices compared to the classical simulation due to noise, particularly affecting the success rate of vertex questions (Fig. 4(b), 6). We give two reasons for this:
  1. 1.
[…]
  2. 2.
    As the strategy depends on entanglement, error in the entangling CX gates disrupts the strategy.
This sensitivity suggests that the win rate of the strategy for $G_{14}$ can evaluate the ability of quantum devices to accurately perform nonlocal operations and serve as a benchmark. In particular, the vertex question win rate is very sensitive to noise, measuring the fidelity of the device gates, whereas the edge question win rate can confirm if a device is using quantum resources. The optimal classical strategy of 4 colors consists of a 4-coloring of $G_{13}$ and assigning the most infrequently used color to the apex vertex $\alpha$. Therefore, all vertex questions would be correctly answered, and 1 edge would be incorrectly answered, resulting in an edge win rate of $36/37\approx 97.2\%$. Any win rate higher than this requires quantum resources. In our experiments, no device exceeded this threshold (Fig. 6). Introducing an error-correcting version of our quantum strategy could improve the robustness of this test, which we leave to future investigation.
4 Conclusion
We present a variational algorithm to compute novel quantum strategies for nonlocal games by encoding the rules of a nonlocal games into a Hamiltonian and using a dual-phase optimization procedure. Our key insight is to optimize separately a state preparation circuit and the measurement scheme, while employing sound circuit initialization and generation techniques, like ADAPT, during optimization. The proposed algorithm is able to recreate many known quantum strategies and also managed to find new short-depth, perfect quantum strategies for a graph on $14$ vertices with no known explicit construction using four qubits. This shows that variational techniques can be used on classical computers to find optimal strategies for small examples of quantum non-local games in cases where analytic techniques fail.
These techniques can also be applied in a quantum setting, wherein sampling is used to estimate the gradient. We analyze the number of operator exponentials needed for such a protocol and provide conditions under which it is efficient as a function of the minimum gradient, the number of training epochs needed to converge, the number of players present, and the learning rate used. However, it is worth noting that the present of barren plateaus is expected with the training objective function, generically meaning that “warm starts” or other techniques to avoid vanishing gradients may be needed to successfully deploy these methods on quantum computers.
We further illustrate these methods as a benchmark on IBM quantum devices, we noticed a significant inability to reproduce our newly found strategy hardware, primarily due to noise affecting the circuit. We believe that graph nonlocal games serve as an excellent benchmark problem for quantum hardware since they allow researchers to evaluate the capabilities of quantum devices as problems scale and hardware architectures become more complex. By evaluating the win rates of both vertex and edge questions in these games, the win rate of vertex questions can gauge the ability of a device to perform nonlocal operations and measure gate fidelity, while the win rate of edge questions can help confirm the utilization of entanglement across a device. Furthermore, our research lays the groundwork for further exploration of nonlocal games strategies that were previously out of reach using the tools and techniques from quantum machine learning.
Data Availability
The code used to generate the data and figures in this article can be found at
https://github.com/jfurches/nonlocalgames.
Acknowledgements
Thanks to David Roberson for providing valuable feedback. JF and COM were supported by the Laboratory Directed Research and Development Program and Mathematics for Artificial Reasoning for Scientific Discovery investment at the Pacific Northwest National Laboratory, a multiprogram national laboratory operated by Battelle for the U.S. Department of Energy under Contract DEAC05- 76RLO1830. NW and COM were funded by grants from the US Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-Design Center for Quantum Advantage under contract number DE-SC0012704.
Appendix A Measurement Parameters
Alice’s measurement parameters of the $G_{14}$ strategy are contained within
data/g14_constrained_u3ry/g14_state.json with the key phi. Constructing this into a NumPy array should return a tensor of shape $(1,14,2,4)$, corresponding to (players, questions, qubits, parameters). This tensor can be transformed to produce the conjugated measurement angles for Bob, as seen in U3RyLayer in measurement.py.
Appendix B Hyperparameters
Problem Hyperparameter Value
CHSH ADAPT Grad Max $\epsilon_{\theta}$ $10^{-3}$
BFGS Grad Max $\epsilon_{\phi}$ $10^{-5}$
DPO Tolerance $\Delta E$ $10^{-3}$
NPS Same as CHSH
$G_{14}$ ADAPT Grad Max $\epsilon_{\theta}$ $10^{-6}$
BFGS Grad Max $\epsilon_{\phi}$ $10^{-5}$
DPO Tolerance $\Delta E$ $10^{-6}$
Classical Circuit Shots $10^{4}$
Quantum Device Shots $1024$
Table 1: Hyperparameters for DPO experiments
We give the algorithm hyperparameters for our experiments. The parameter $\epsilon_{\theta}$ refers to the convergence criteria of ADAPT used to prepare the shared state |$\ket{\psi(\theta)}$. ADAPT finishes when the maximum pool gradient element reaches the threshold, $\max_{A_{i}}\left|\braket{[H,A_{i}]}\right|<\epsilon_{\theta}$. Similarly, the parameter $\epsilon_{\phi}$ controls the convergence of the second phase of DPO, as the BFGS optimizer halts when $\max_{i}\left|\nabla_{\phi}\braket{H}\right|<\epsilon_{\phi}$. Finally, $\Delta E$ controls the termination of the overall DPO procedure, ending when $\braket{H^{(k-1)}}-\braket{H^{(k)}}<\Delta E$ at iteration $k$.
Appendix C Quantum Strategy as a Database
In this section, we view the quantum strategy shared by the players of a non-local game as access to a shared quantum database, and illustrate how entanglement provides an advantage in non-local games. As before, we denote the measurement operators as $\mathcal{M}_{a|q}=U_{q}^{\dagger}P_{a}U_{q}$. Since upon each question, the players perform the transformation |$\ket{\psi}\rightarrow U_{q}\ket{\psi}$, which can be viewed as selecting a table from |$\ket{\psi}$ (the database) based on the classical question $q$.
The unitary $U_{q}$ maps from |$\ket{\psi}\rightarrow\ket{a}$ and can be decomposed into $U_{q}=\sum_{j}\ket{k_{j}}\bra{e_{j}}$ |, where |$\ket{e_{j}}$ are orthonormal vectors, and |$\ket{k_{j}}$ are considered keys in the space of |$\ket{a}$. Using this, we obtain
[…]
and the overlap between the answer and state is given by
[…]
$=\sum_{i}\braket{e_{iq}}{\psi}\braket{a}{k_{iq}},$ (19)
with the condition that $\sum_{i}\left|\braket{e_{iq}}{\psi}\right|^{2}=1$.
The attention mechanism in machine learning [26], a neural differentiable database, is given by
$\text{attention}(q,k,v)=\text{softmax}\left(\frac{qk^{T}}{\sqrt{d_{k}}}\right)v,$ (20)
where $q$ is a matrix of query vectors, $k$ is a matrix of key vectors, $v$ is a value matrix, and $d_{k}$ is a normalization factor of the key matrix dimension. The function $\text{softmax}(\cdot)$ defines a probability distribution to weight $v$ by the similarity of $q$ and $k$. Analogously, one can view $\braket{e_{iq}}{\psi}$ as a complex attention weight that scales the values $\braket{a}{k_{iq}}\in\mathbb{C}$. This is illustrated in fig. 7.
Figure 7: Database structure of strategy.
We show that entanglement gives more control over the weights compared to a product state, providing a more expressive strategy. First, we start with a product state
|$\ket{\psi}$ $=\bigotimes_{i}\ket{\psi_{i}}$
$=\left(\sum_{j}c_{j}\ket{j}\right)\otimes\left(\sum_{l}d_{l}\ket{l}\right)$
$=\sum_{jl}c_{j}d_{l}\ket{jl}$
where |$\ket{j},\ket{l}$ are basis vectors. Computing the overlap $\braket{a}{U_{q}}{\psi}$, we obtain
$\braket{a}{U_{q}}{\psi}$ $=\sum_{i}\braket{a}{k_{i}}\sum_{jl}c_{j}d_{l}\braket{e_{i}}{jl}$ (21)
$=\sum_{i}w^{c}_{i}v_{i},$ (22)
where $v_{i}=\braket{a}{k_{i}}$ and $i=1,\dots,2^{N_{a}+N_{b}}$ enumerates the number of basis vectors of $U_{q}$. We denote the classical attention weights by $w_{i}^{c}=\sum_{jl}c_{j}d_{l}\braket{e_{i}}{jl}$. Similarly, for an entangled state |$\ket{\psi}=\sum_{jl}c_{jl}\ket{jl}$, we obtain
$\braket{a}{U_{q}}{\psi}$ $=\sum_{i}\braket{a}{k_{i}}\sum_{jl}c_{jl}\braket{e_{i}}{jl}$ (23)
$=\sum_{i}w_{i}^{q}v_{i},$ (24)
where the quantum weights are $w_{i}^{q}=\sum_{jl}c_{jl}\braket{e_{i}}{jl}$. Comparing $w_{i}^{c}$ and $w_{i}^{q}$, the classical weights have $|j|+|l|-2=\mathcal{O}(2^{N_{a}}+2^{N_{b}})$ free parameters (controllable through choice of shared state), while the quantum weights have $|j||l|-1=\mathcal{O}(2^{N_{a}+N_{b}})$ free parameters, where the $-1$ and $-2$ terms come from normalizing the quantum states. The larger parameter count results in much more granular control over the resulting distribution $p(\mathbf{a}|\mathbf{q})=|\mathbf{w}^{T}\mathbf{v}|^{2}$, giving entangled shared states an advantage in the non-local game.